a)



An increase in the ionisation of a molecule in a GI environment will usually



increase its solubility. Valproic acid is an oral dosage form used in epilepsy and



has a pKa of 4.8. With the stomach pH at 2.0 and the ileum at pH 6.8, which area



of the GIT will Valproic acid be more soluble.






b)



You are required to manufacture 1L of 50mM pH 7.0 phosphate buffer that could be



used for an injectable formulation of valproic acid. The Ka of dihydrogen phosphate is



6.3x10


-8


and the RMM of potassium dihydrogen phosphate and dipotassium



hydrogen phosphate are 136 and 174 respectively.



I.



Calculate the pKa of dihydrogen phosphate.



II.



What is the ratio between the conjugate base and the acid to be used?



III.



Calculate the mass of anhydrous potassium dihydrogen phosphate and



dipotassium hydrogen phosphate with RMM of 136 and 174, and



concentrations of 30.5 and 19.5 mmol respectively.




1
Expert's answer
2022-03-14T21:42:05-0400

a) pH = pKa + log10{[A-]/[HA]}

Stomach (pH = 2.0)

2.0 = 4.8 + log10{[A-]/[HA]}

-2.8 = log10{[A-]/[HA]}

2.8 = log10{[HA]/[A-]}

631x + x = [A]total

99.84% is protonated and 0.16% is unprotonated

Ileum (pH = 6.8)

6.8 = 4.8 + log10{[A-]/[HA]}

2.0 = log10{[A-]/[HA]}

100x + x = [A]total

99.00% is unprotonated and 1.00% is protonated

In stomach Valproic acid is more soluble

b) I. pKa = -log10 Ka = -log10 (6.3×10-8) = 7.2

II. pH = pKa + log{[A-]/[HA]}

[HA] + [A-] = 0.05

7.0 = 7.2 + log{(0.05-x)/x}

x/(0.05 - x) = 1.58

x = 0.0306

[HA] = 0.0306 mol

[A-] = 0.0194 mol

[A-]/[HA] = 1.577

III. For molecular weights: KH2PO4 = 39 + 97 = 136 amu; K2HPO4 = 39×2 + 96 = 174 amu

Thus, 0.0306×136 = 4.161 g of KH2PO4 and 0.0194×174 = 3.376 g of K2HPO4 should be added.